A CD player is on sale for $270. What was the original price of the CD player if the discount was 2/5 of the original price?
Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

270/.060=$450

Answer by mananth(16949)

let x be the original price
Discount is 2/5 of x
sale price = 270
..
x-2x/5 = 270
(5x-2x)/5 = 270
3x/5 =270
3x=1350
/3
x= 450
= $ 450 original price
...
m.ananth@hotmail.ca